Please help me to solve this problem.I really don`t understand and drastically need help. Thank you so much for all that you do! Please help me! Here is the problem:

Tommy has 40 coins(all nickels, dimes, and quarters) worth 4.50. He has 7 more nickels than dimes. How many quarters does Tommy have? Write an equation.

Thanks again so much for the help!

Answer by stanbon(48535)

Tommy has 40 coins(all nickels, dimes, and quarters) worth 4.50. He has 7 more nickels than dimes. How many quarters does Tommy have? Write an equation.
-------------------------
Let # of dimes be "x"; Value of these is 10x cents
Number of nickels is "x+7" ; Value of these is 5(x+7) = 5x+35 cents
Number of quarters is 40 - x - (x+7)= 33-2x ; Value = 25(33-2x)= 825-50x cents
-----------------------------------------------
EQUATION:
value + value + value = 450 cents
10x + 5x + 35 + 825 - 50x = 450
-35x + 860 = 450
-35x = -410
x = 11.71
Note: The solution should be a whole number as it
is the number of dimes.
Check your problem statement.
=====================================
Cheers,
Stan H.